5. An 8.0 mm (0.31 in.) diameter cylindrical rod fabricated from a 7075-T6 alloy is subjected to reversed tension-compressionCan you show step by step how to solve it, including the equations.
Answer #1

Provided Informations in problem are as below

Cylindrical rod diameter = 8mm

Thus area A= π*82/4= 50.27 mm​​​​​2

Material used 7076-T6 Alloy which is having following properties ( from material standard)

Ultimate tensile strength = S​​​​​ut​= 572 MPa

Tensile yield strength = S​​​​​Sy = 503 Mpa

Fatigue strength = S​​​​​​e​​​= 159 MPa

Maximum tensile L​​​​​​t and Lc Compressive loads= +8400 N and -8400 N

Thus maximum tensile(S​​​​​​t) and compressive (Sc)stresses = +(8400/50.27) and -(8400/50.27) = +167.1 N/mm​​​2​​​​ and - 167.1 N/mm​​​​​2

Find out fatigue life

Equations for fatigue life are as below

N = 10-c/b*S​​​​​​a​​​​​1/b

b=( -1/3)* log (0.8*S​​​​​​ut/S​​​​​e​​​) = (-1/3)*log [(0.8*572/159)]= - 0.15393

c= log [(0.8*S​​​​​​ut)2/S​​​​​​e​​​] = log [(0.8*572)2/159] = 3.12

Where N = Fatigue life in no of cycles

S​​​​​​a​​​= Stress amplitude= [S​​​​​t-(-S​​​​​c​)]/2 = 167.1 N/ mm​​2

Substituting known values in above equation we can compute Fatigue life as below.

N= 10-(3.12/(-0.15393))*167.21/(-0.15393)

= 1020.79*167.2-6.5= 2182550 Cycles